OpenStudy (anonymous):

Calculus: Evaluate f(x)=1/x at the specified value of the independent variable and simplify the result. (f(1+Δx)-f(1))/Δx

OpenStudy (anonymous):

\[f(x)=1/x\] \[\frac{ f(1+Δx)-f(1) }{ Δx }\]

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

@DebbieG

OpenStudy (anonymous):

Um... does anyone know to how to solve this?

OpenStudy (debbieg):

So is this a differential?

OpenStudy (anonymous):

I just started my calculus class, so we're reviewing with substitutions

OpenStudy (debbieg):

Dont you just treat the Δx like we do the "h" in the difference quotient? (I'm not terribly familiar with differentials... but isn't it just a riff on the difference quotient?)

OpenStudy (debbieg):

ok.. well... my best guess is to treat is like a diff quotient. Just use that Δx in place of h. So you need to use the function: \(\Large \dfrac{ f(1+Δx)-f(1) }{ Δx }=\dfrac{\dfrac{1}{1+Δx}-1 }{ Δx }\) and then simplify that....

OpenStudy (debbieg):

unless you know what Δx is...if you were given another x value somewhere?

OpenStudy (anonymous):

no the problem is as I wrote it

OpenStudy (debbieg):

Differentials are weird. I guess it would go something like this?.... (but I'm not terribly confident): \(\Large \left(\dfrac{1}{1+Δx}-1 \right)\cdot\dfrac{1}{ Δx }\\ \Large \left(\dfrac{1-(1+Δx)}{1+Δx} \right)\cdot\dfrac{1}{ Δx }\\ \Large \left(\dfrac{-Δx}{1+Δx} \right)\cdot\dfrac{1}{ Δx }\\ \Large \dfrac{-1}{1+Δx} \)

OpenStudy (debbieg):

but it kind of makes sense, since as Δx->0 that expression goes to -1, and the derivative here \(f\prime(1)=-1\).

OpenStudy (anonymous):

I don't understand how you went from step 1 to step 2

OpenStudy (debbieg):

@southpaw, I'm sorry, I had to run out unexpectedly.... All I did from step 1 to step 2 was to join everything inside the () over the common den'r.... \(\Large \Large \left(\dfrac{1}{1+Δx}-1 \right)\cdot\dfrac{1}{ Δx }\\ \Large \left(\dfrac{1}{1+Δx}-\dfrac{1+Δx}{1+Δx} \right)\cdot\dfrac{1}{ Δx }\\ \Large \left(\dfrac{1-(1+Δx)}{1+Δx} \right)\cdot\dfrac{1}{ Δx }\\ \Large \left(\dfrac{-Δx}{1+Δx} \right)\cdot\dfrac{1}{ Δx }\\ \Large \dfrac{-1}{1+Δx}\)

OpenStudy (anonymous):

alright thank you